Find the sales tax and total cost of an item that cost $109.99 if the tax rate is 7%
Found 2 solutions by Maths68, Fals:
Answer by Maths68(1474)

Cost = $109.99
Sales Tax = 109.99 * 7% = 109.99 * 7/100 = 1.0999 * 7 = 7.6993 = $7.70
Total Cost = Sales Tax + Cost = 7.70+109.99= $117.69

Answer by Fals(5)

109.99 times .07 = 7.70
$7.70 is the sales tax for this item.
109.99+7.70=117.69
$117.69 is the total price for the item.